# AttnLRP: Attention-Aware Layer-wise Relevance Propagation for Transformers

### v1

@@ B.5 Attributions on SQuAD v2
Gradient-based methods such as G×$\times$×I, SmoothGrad, IG or Grad-CAM are noisy and not really informative. While the heatmaps of I×$\times$×G or Grad-CAM seem to be inverted, we experimented with inverting the attributions on a subset, however we did not notice improvement and applied the rules with their original definition. Grad×$\times$×Attn Rollout suffers from background noise. AtMan produces highly sparse attributions, assigning large positive relevance to the answer token 18, however, also assigning a similar amount of relevance to the token much, which is part of the question. AttnLRP and CP-LRP identify the token 18 as being the most relevant token and also relate it (by assigning positive and negative relevance) to other information in the text such as 27.7, 132 or average. We conjecture that such targeted contrasting reflects the reasoning process of the model (e.g., is necessary to distinguish between related questions about how many tons are blown out vs. how many tons remain on the ground). A systematic analysis of these effects remain an interesting topic for future work. The similarity between AttnLRP and CP-LRP in Flan-T5 are in line with the quantitative evaluation from Table 1, which showed a small, but consistent advantage of AttnLRP over CP-LRP wrt. faithfulness, while in Llama 2, AttnLRP substantially outperforms. For comparison, we also visualize a random attribution with Gaussian noise.
@@
Figure 7: Evaluation on the SQuAD v2: We compute attributions for different state-of-the-art methods on the first token of the answer (highlighted in red). Gradient-based methods such as G×$\times$×I, SmoothGrad, IG or Grad-CAM are noisy. Grad×$\times$×Attn Rollout suffers from background noise. AtMan produces highly sparse attributions, assigning an equal amount of relevance to a token, which is part of the question, as to token 18. CP-LRP has a different weighting of the tokens e.g. the word ‘much’ in the question is not highlighted by CP-LRP, while AttnLRP highlights it stronger and AtMan focuses excessively on it. For comparison, we also visualize a random attribution with Gaussian noise.
B.6 Benchmarking Cost, Time and Memory Consumption
@@
To calculate energy cost, we assume a price of $0.16$ per kWh of energy, and that a single A100 GPU consumes on average 130W. Figure 8 depicts the cost, the runtime and peak GPU memory consumption. Since perturbation-based methods are memory efficient, a 70b model with full context size of 4096 is attributable. However, LRP with checkpointing requires more memory than a node can supply.
Figure 8: From left to right: Cost in dollar, time in seconds and peak GPU memory in gigabytes for AttnLRP and linear-time perturbation. Evaluated on Llama 2-70b and Llama 2-13b models on a node with four A100-SXM4 40GB. G×$\times$×AttnRollout is in the range of AttnLRP and omitted for clarity of visualization. Because AttnLRP consumes more than 160 GB of RAM, the curves for the 70b model stop. Measured at fixed intervals of context size 32, 64, 128, 256, 512, 1024, 2048, 3000, 4096.
B.7 Attributions of Knowledge Neurons
Figure 9, 10 and 11 illustrate the top 10 sentences in the Wikipedia summary dataset that maximally activate a knowledge neuron. We applied AttnLRP to highlight the tokens inside these reference samples. We observe that knowledge neurons exhibit remarkable disentanglement, e.g., neuron #256 of layer 18 shown in Figure 9 seems to encode concepts related to transport systems (railways in particular), while neuron #2207 of layer 20 shown in Figure 10 seems to encode the concept teacher, in particular a teacher, in an unusual context (e.g., inappropriate behavior, sexual misconduct). The degree of disentanglement should be studied in future work.
Figure 9: AttnLRP attributions on top 10 ActMax sentences collected over the Wikipedia summary dataset for neuron #256, in layer 18. The knowledge neuron seems to activate for transport systems (railways in particular).
Figure 10: AttnLRP attributions on top 10 ActMax sentences collected over the Wikipedia summary dataset for neuron #2207, in layer 20. The knowledge neuron is activating for ‘teacher’, in unusual context such as inappropriate behavior, sexual misconduct etc.
Figure 11: AttnLRP attributions on top 10 ActMax sentences collected over the Wikipedia summary dataset for neuron #922, in layer 18. The knowledge neuron seems to be activating for scientific descriptions of plants.
Table 5: ViT Perturbation Experiment (Faithfulness). For SmoothGrad, we set $\sigma=0.01$, for AtMan $p=1.0$ and $t=0.1$, for AttnRoll $dt=0.99$, and for G×$\times$×AttnRoll $dt=0.91$. “all epsilon” indicates that the $\varepsilon$-rule has been used on the linear and convolutional layers. The term “best” refers to the utilization of LRP with the composite proposed in B.4. $\Delta A^{F}$ denotes the area under the curve for a flipping perturbation experiment which leverages both $A_{MoRF}^{F}$ of the most relevant first order, and $A_{LeRF}^{F}$ of least relevant first order. ($\Delta A^{F}=A_{LeRF}^{F}-A_{MoRF}^{F}$). As discussed in Section B.2, this is equivalent to insertion perturbation.
@@ B.7 Attributions of Knowledge Neurons
Table 6: Wikipedia Perturbation Experiment (Faithfulness). For SmoothGrad, we set $\sigma=0.1$, for AtMan $p=1.0$, for AttnRoll $dt=1$, and for G×$\times$×AttnRoll $dt=1$. ”all epsilon” indicates that the $\varepsilon$-rule has been used on all linear layers. $\Delta A^{I}$ denotes the area under the curve for the insertion perturbation experiment which leverages both $A_{MoRF}^{I}$ of the most relevant first order, and $A_{LeRF}^{I}$ of least relevant first order. ($\Delta A^{I}=A_{MoRF}^{I}-A_{LeRF}^{I}$). As discussed in Section B.2, this is equivalent to flipping perturbation.
@@
Table 7: IMDB Perturbation Experiment (Faithfulness), For SmoothGrad we set $\sigma=0.05$, for AtMan $p=0.7$, for AttnRoll $dt=1$, and for G×$\times$×AttnRoll $dt=1$. ”all epsilon” indicates that the $\varepsilon$-rule has been used to propagate relevance to the layers. $\Delta A^{I}$ demonstrates the area under the curve for the perturbation experiment of the type Insertion which leverages insights from both $A_{MoRF}^{I}$ of the most relevant first order, and $A_{LeRF}^{I}$ of least relevant first order. ($\Delta A^{I}=A_{MoRF}^{I}-A_{LeRF}^{I}$). As discussed in Section B.2, this is equivalent to flipping perturbation.
@@
Table 8: SQuAD v2 Perturbation Experiment (Faithfulness). For SmoothGrad, we set $\sigma=0.1$ and for AtMan $p=0.9$, for AttnRoll $dt=1$, and for G×$\times$×AttnRoll $dt=1$. ”all epsilon” indicates that the $\varepsilon$-rule has been used on all linear layers. $\Delta A^{I}$ denotes the area under the curve for the insertion perturbation experiment which leverages both $A_{MoRF}^{I}$ of the most relevant first order, and $A_{LeRF}^{I}$ of least relevant first order. ($\Delta A^{I}=A_{MoRF}^{I}-A_{LeRF}^{I}$). As discussed in Section B.2, this is equivalent to flipping perturbation.
Methods Flan-T5-XL
SQuAD v2
($\uparrow$)$\Delta A^{I}$ ($\uparrow$)$A_{MoRF}^{I}$ ($\downarrow$)$A_{LeRF}^{I}$
Random 0.01 -6.05 -6.06
I×$\times$×G 0.27 -5.51 -5.78
IG 0.77 -5.31 -6.08
SmoothG 0.16 -5.47 -5.63
GradCAM 0.94 -5.24 -6.19
AttnRoll -0.42 -5.96 -5.54
G×$\times$×AttnRoll -0.06 -5.65 -5.60
AtMan 1.01 -5.18 -6.19
CP-LRP (all epsilon) 1.74 -4.93 -6.67
AttnLRP (all epsilon) 1.76 -4.91 -6.67
Figure 12: Statistics on Rules used for softmax layers: Either applying $z^{+}$, $\varepsilon$-rule, or regarding as constant as proposed in CP-LRP. Propagating relevance values through (specifically by applying $z^{+}$ rule) softmax improves the faithfulness of explanations compared to the case where we block its propagation.
Figure 13: Statistics on Rules used for Convolution layers: Applying $z^{+}$ and AlphaBeta proposes acceptable results however the most faithful results can be reached via Gamma($\gamma=0.25$).
Figure 14: Statistics on Rules used for Linear layers: Similar to Convolution layers, Gamma seams more promising however with different $\gamma$ value ($0.05$ in this case).
Figure 15: Statistics on Rules used for LinearInputProjection layers: Gamma and $\epsilon$ rules are competitive in this case, however since there is larger difference between the minimum and the lower quartile in Gamma rules, the most faithful choice will be $\epsilon$-rule.
Figure 16: Statistics on Rules used for LinearOutputProjection layers: The $\epsilon$-rule outperforms other rules clearly.

### v2

@@ B.7 Attributions on SQuAD v2
In Figure B.9, the heatmaps of I×$\times$×G or Grad-CAM seem to be inverted, hence we experimented with inverting the attributions on a subset, however we did not notice improvement and applied the rules with their original definition. AtMan produces highly sparse attributions, assigning large positive relevance to the answer token 18, however, also assigning a similar amount of relevance to the token much, which is part of the question. AttnLRP and CP-LRP identify the token 18 as being the most relevant token and also relate it (by assigning positive and negative relevance) to other information in the text such as 27.7, 132 or average. We conjecture that such targeted contrasting reflects the reasoning process of the model (e.g., is necessary to distinguish between related questions about how many tons are blown out vs. how many tons remain on the ground). A systematic analysis of these effects remain an interesting topic for future work.
@@
Figure B.7: Evaluation on the Mixtral 8x7b model: We compute attributions for different state-of-the-art methods for the answer token “France”. Gradient-based methods such as G×$\times$×I, SmoothGrad, IG or Grad-CAM are noisy. Grad×$\times$×Attn Rollout suffers from background noise. While AtMan usually generates sparse heatmap, in this case it fails (compare Figure B.9). CP-LRP highlights “Normandy” the strongest, while AttnLRP highlights the correct token “France”. For comparison, we also visualize a random attribution with Gaussian noise.
Figure B.8: Evaluation on the Mixtral 8x7b model: We compute attributions for different state-of-the-art methods for all tokens at the same time inside the answer “Ibn Sina”. Gradient-based methods such as G×$\times$×I, SmoothGrad and IG are noisy. Grad-CAM highlights the correct tokens except it misses the beginning token “I” of the word “Ibn”. Likewise AtMan fails to highlight all tokens. Grad×$\times$×Attn Rollout suffers from background noise. CP-LRP resembles random noise, while AttnLRP highlights the correct tokens “Ibn Sina” in its entirety. For comparison, we also visualize a random attribution with Gaussian noise.
Figure B.9: Evaluation on the Flan-T5-XL model: We compute attributions for different state-of-the-art methods on the first token of the answer (highlighted in red). Gradient-based methods such as G×$\times$×I, SmoothGrad, IG or Grad-CAM are noisy. Grad×$\times$×Attn Rollout suffers from background noise. AtMan produces highly sparse attributions, assigning an equal amount of relevance to a token, which is part of the question, as to token 18. CP-LRP has a different weighting of the tokens e.g. the word ‘much’ in the question is not highlighted by CP-LRP, while AttnLRP highlights it stronger and AtMan focuses excessively on it. For comparison, we also visualize a random attribution with Gaussian noise.
B.8 Benchmarking Cost, Time and Memory Consumption
@@
To calculate energy cost, we assume a price of $0.16$ per kWh of energy, and that a single A100 GPU consumes on average 130W. Figure B.10 depicts the cost, the runtime and peak GPU memory consumption. Since perturbation-based methods are memory efficient, a 70b model with full context size of 4096 is attributable. However, LRP with checkpointing requires more memory than a node supplies.
Figure B.10: From left to right: Cost in dollar, time in seconds and peak GPU memory in gigabytes for AttnLRP and linear-time perturbation. Evaluated on LLaMa 2-70b and LLaMa 2-13b models on a node with four A100-SXM4 40GB. G×$\times$×AttnRollout is in the range of AttnLRP and omitted for clarity of visualization. Because AttnLRP consumes more than 160 GB of RAM, the curves for the 70b model stop. Measured at fixed intervals of context size 32, 64, 128, 256, 512, 1024, 2048, 3000, 4096.
B.9 Attributions of Knowledge Neurons
Figure B.11, B.12 and B.13 illustrate the top 10 sentences in the Wikipedia summary dataset that maximally activate a knowledge neuron. We applied AttnLRP to highlight the tokens inside these reference samples. We observe that knowledge neurons exhibit remarkable disentanglement, e.g., neuron #256 of layer 18 shown in Figure B.11 seems to encode concepts related to transport systems (railways in particular), while neuron #2207 of layer 20 shown in Figure B.12 seems to encode the concept teacher, in particular a teacher, in an unusual context (e.g., inappropriate behavior, sexual misconduct). The degree of disentanglement should be studied in future work.
Figure B.11: AttnLRP attributions on top 10 ActMax sentences collected over the Wikipedia summary dataset for neuron #256, in layer 18. The knowledge neuron seems to activate for transport systems (railways in particular).
Figure B.12: AttnLRP attributions on top 10 ActMax sentences collected over the Wikipedia summary dataset for neuron #2207, in layer 20. The knowledge neuron is activating for ‘teacher’, in unusual context such as inappropriate behavior, sexual misconduct etc.
Figure B.13: AttnLRP attributions on top 10 ActMax sentences collected over the Wikipedia summary dataset for neuron #922, in layer 18. The knowledge neuron seems to be activating for scientific descriptions of plants.
Table B.7: ViT-B-16 Perturbation Experiment (Faithfulness). For SmoothGrad, we set $\sigma=0.01$, for AtMan $p=1.0$ and $t=0.1$, for AttnRoll $dt=0.99$, and for G×$\times$×AttnRoll $dt=0.91$. “all epsilon” indicates that the $\varepsilon$-rule has been used on the linear and convolutional layers. The term “best” refers to the utilization of LRP with the composite proposed in B.5. $\Delta A^{F}$ denotes the area under the curve for a flipping perturbation experiment which leverages both $A_{MoRF}^{F}$ of the most relevant first order, and $A_{LeRF}^{F}$ of least relevant first order. ($\Delta A^{F}=A_{LeRF}^{F}-A_{MoRF}^{F}$). As discussed in Section B.2, this is equivalent to insertion perturbation.
@@ B.9 Attributions of Knowledge Neurons
Table B.8: Wikipedia Perturbation Experiment (Faithfulness). For SmoothGrad, we set $\sigma=0.1$, for AtMan $p=1.0$, for AttnRoll $dt=1$, and for G×$\times$×AttnRoll $dt=1$. ”all epsilon” indicates that the $\varepsilon$-rule has been used on all linear layers. $\Delta A^{I}$ denotes the area under the curve for the insertion perturbation experiment which leverages both $A_{MoRF}^{I}$ of the most relevant first order, and $A_{LeRF}^{I}$ of least relevant first order. ($\Delta A^{I}=A_{MoRF}^{I}-A_{LeRF}^{I}$). As discussed in Section B.2, this is equivalent to flipping perturbation.
@@
Table B.9: IMDB Perturbation Experiment (Faithfulness), For SmoothGrad we set $\sigma=0.05$, for AtMan $p=0.7$, for AttnRoll $dt=1$, and for G×$\times$×AttnRoll $dt=1$. ”all epsilon” indicates that the $\varepsilon$-rule has been used to propagate relevance to the layers. $\Delta A^{I}$ demonstrates the area under the curve for the perturbation experiment of the type Insertion which leverages insights from both $A_{MoRF}^{I}$ of the most relevant first order, and $A_{LeRF}^{I}$ of least relevant first order. ($\Delta A^{I}=A_{MoRF}^{I}-A_{LeRF}^{I}$). As discussed in Section B.2, this is equivalent to flipping perturbation.
@@
Figure B.14: Statistics on Rules used for softmax layers: Either applying $z^{+}$, $\varepsilon$-rule, or regarding as constant as proposed in CP-LRP. Propagating relevance values through (specifically by applying $z^{+}$ rule) softmax improves the faithfulness of explanations compared to the case where we block its propagation.
Figure B.15: Statistics on Rules used for Convolution layers: Applying $z^{+}$ and AlphaBeta proposes acceptable results however the most faithful results can be reached via Gamma($\gamma=0.25$).
Figure B.16: Statistics on Rules used for Linear layers: Similar to Convolution layers, Gamma seems more promising however with different $\gamma$ value ($0.05$ in this case).
Figure B.17: Statistics on Rules used for LinearInputProjection layers: Gamma and $\epsilon$ rules are competitive in this case, however since there is larger difference between the minimum and the lower quartile in Gamma rules, the most faithful choice will be $\epsilon$-rule.
Figure B.18: Statistics on Rules used for LinearOutputProjection layers: The $\epsilon$-rule outperforms other rules clearly.